OpenStudy (anonymous):

use the binomial theorem to find the binomial expansion of expression (x-3)^5

OpenStudy (campbell_st):

well binomial theorem says \[(a+b)^n = \sum_{k=0}^{n} \left(\begin{matrix}n \\ k\end{matrix}\right) (a)^{n - k} (b)^k\] so you have a = x, b = -3 and n = 5 you can substitute them and calculate